OpenStudy (anonymous):

determine the coordinates of a point on the line with the symmetric equation (x+3)/2=(y-2)/4=(z+4)/3

OpenStudy (anonymous):

\[\frac{ x+3 }{ 2 } = \frac{ y-2 }{ 4 } = \frac{ z+4 }{ 3 }\]

OpenStudy (jtvatsim):

These are fun. All you need to do is pick values for x, y, and z that make all sides equal to each other.

OpenStudy (jtvatsim):

We can begin with any of the "fractions" so how about we pick something for x, see what fraction we get, and then try to pick y and z to match.

OpenStudy (jtvatsim):

So, what number would you like to pick for x?

OpenStudy (anonymous):

would i need to have common denominators?

OpenStudy (jtvatsim):

Not necessarily, but it isn't a bad idea if you want to go that way.

OpenStudy (jtvatsim):

It would certainly make guessing a lot easier.

OpenStudy (jtvatsim):

Let's try that and see.

OpenStudy (anonymous):

ok so \[\frac{ 6x+18 }{ 12 } = \frac{ 3y-6 }{ 12 } = \frac{ 4z+16 }{ 12 }\]

OpenStudy (jtvatsim):

Alright, looking good.

OpenStudy (jtvatsim):

Any ideas on how you want to move forward?

OpenStudy (anonymous):

im not too sure. would I combine them to (6x+18+3y-6+4z+16)/12 and then simplify?

OpenStudy (jtvatsim):

Well, we can't combine them quite like that since we don't know exactly what they equal. But we do know that they are all equal to the same thing. Let's call the thing they equal t. So, \[\frac{6x + 18}{12} = \frac{3y-6}{12} = \frac{4z + 16}{12} = t\] Are you ok with that so far?

OpenStudy (anonymous):

yes i am

OpenStudy (jtvatsim):

Alright, so what we can do is solve each equation for x, y, and z. I'll do the one for x. \[\frac{6x + 18}{12} = t\] so \[x= (12t -18)/6 = 2t - 3\]

OpenStudy (jtvatsim):

Do you see how I did that?

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

would i do the same for y and z?

OpenStudy (jtvatsim):

Yes do the same for y and z.

OpenStudy (anonymous):

y=(12t+6)/3 z=(12-16)/4

OpenStudy (jtvatsim):

I think a t might have escaped for the z, double check that one. :)

OpenStudy (jtvatsim):

Your y is correct though.

OpenStudy (jtvatsim):

Oh I see, I think you meant, z = (12t - 16)/4 right?

OpenStudy (anonymous):

oh yes sry about that

OpenStudy (jtvatsim):

So we can simplify our answers and we see that x = 2t - 3 y = 4t + 2 z = 3t - 4 Does that make sense so far, we are nearly done!

OpenStudy (anonymous):

yes i got so far

OpenStudy (jtvatsim):

OK, so now we have the equations to tell us all of the points on the line! Just plug in any t you want and our equations will give you the x,y,z coordinate. For example, For t = 0, x = -3 y = 2 z = -4 The point (-3, 2, -4) is on the line.

OpenStudy (anonymous):

ok cool! thanks a lot!

OpenStudy (jtvatsim):

Or you could plug in t = 1: x = -1, y = 6, z = -1 is also a point on the line!

OpenStudy (jtvatsim):

Quick shortcut for you.

OpenStudy (jtvatsim):

We actually didn't need to find common denominators!

OpenStudy (jtvatsim):

Notice that we could just take the original fractions and set them equal to t.

OpenStudy (anonymous):

haha just realized!

OpenStudy (jtvatsim):

This gives us the same equations! But I wanted to give you the chance to be in the "driver's seat" and find the solution in your own way. Good work!

OpenStudy (anonymous):

thanks!